use the method of Frobenius to Solve xy+2y+4 xy=0
Answer #1

Now, nem-1 ys dy da Z Cn (nem) x n=o and Nem-2 dy y = I Cn (nem) (nem-1) dra n=o Now putting the value of y, yandy in (i)+ 4 Cn-2=0 +4 Cn-2= Cn nemt { (nem) => Cn - 4 Cn-2 (*+m) (n=m+1) n72 (vi) equating the co-efficient of the term of the next h— ЧСИ Co (m+ с) (mp7) – Ч Со (m+2) (3) (ms) (m+r) (nv) (m+7) Now from the equation - = Cnx тет Т= 0 mel те). = Com С. x+С, x(m+2) (m+3) 16x4 (m+2) (m+3)(m+4) (m+5) (vii) m = 0 Now putting co by and replacing we get from - (vii) y = a (-***** y putti